# Using Large Multimodal Models to Extract Knowledge Components for Knowledge Tracing from Multimedia Question Information

### v1

@@ Abstract
Knowledge tracing models have enabled a range of intelligent tutoring systems to provide feedback to students. However, existing methods for knowledge tracing in learning sciences are predominantly reliant on statistical data and instructor-defined knowledge components, making it challenging to integrate AI-generated educational content with traditional established methods. We propose a method for automatically extracting knowledge components from educational content using instruction-tuned large multimodal models. We validate this approach by comprehensively evaluating it against knowledge tracing benchmarks in five domains. Our results indicate that the automatically extracted knowledge components can effectively replace human-tagged labels, offering a promising direction for enhancing intelligent tutoring systems in limited-data scenarios, achieving more explainable assessments in educational settings, and laying the groundwork for automated assessment.
Introduction
@@
Intelligent Tutoring Systems (ITS) are advanced computer programs that provide personalized and adaptive educational instruction to learners, and for over half a century they have been a subject of active research and discussion in the interdisciplinary field of education and artificial intelligence (Mousavinasab et al. 2021). These systems integrate techniques from artificial intelligence to deliver tailored instruction, dynamically adjusting to the needs and progress of individual students. By simulating one-on-one tutoring, ITSs offer immediate and specific feedback, enhancing student engagement and improving learning outcomes. This combination of real-time adaptability and personalized feedback makes ITSs a valuable tool in modern education, bridging the gap between traditional classroom methods and individualized learning approaches (Nwana 1990).
@@
Knowledge Tracing (KT) is a foundational task in Intelligent Tutoring Systems (ITS), aiming to model a student’s knowledge state and predict future performance on educational tasks. Traditional KT models rely heavily on statistical techniques to analyze historical problem-solving data. Early approaches, such as Bayesian Knowledge Tracing (BKT), used hidden Markov models to estimate the probability of a student knowing a particular skill at any given time (Corbett and Anderson 1994).
@@
More recent models have leveraged advancements in machine learning, particularly deep learning, to enhance predictive accuracy (Abdelrahman, Wang, and Nunes 2023). Deep Knowledge Tracing (DKT) was one of the first models to apply recurrent neural networks (RNNs) to KT, demonstrating significant improvements over BKT by capturing the sequential nature of students’ learning processes. Subsequent models such as Self-Attentive Knowledge Tracing (SAKT) and Separated Self-AttentIve Neural Knowledge Tracing (SAINT) have further refined these approaches (Choi et al. 2020; Pandey and Karypis 2019).
@@
Knowledge Components (KCs) are defined within the Knowledge-Learning-Instruction framework as acquired units of cognitive function or structure that can be inferred from performance on a set of related tasks (Koedinger, Corbett, and Perfetti 2012). While we cannot directly observe the changes in a student’s KCs, we infer them through interactions during assessment and instruction. In Knowledge Tracing, each problem is often labeled with its corresponding KCs, which are typically assigned by human experts based on their expected relevance to the problem and statistically validated for their ability to explain student performance.
@@
There are several methods to define these KCs. (Pelánek 2017) introduces four categories of domain modeling: ’KCs as disjoint sets’, ’Multiple KCs per item’, ’Hierarchy of KCs’, and ’KCs with prerequisites’. In the KT domain, the most common approaches are modeling KCs as disjoint sets or mapping multiple KCs to an item. The simplest approach is to map each question to a single KC, as used in Bayesian KT, SAKT, and SAINT. When mapping multiple KCs, the relationship between KCs and items is represented in a Q-matrix, where rows represent questions and columns represent KCs. Cognitive Diagnosis Models (CDM) like DINA, NIDA, and generalized DINA use this Q-matrix for KT (De La Torre 2011), thus supporting multiple KCs labeled for each question.
@@
Well-defined KCs are essential for accurately predicting learner performance, building effective adaptive learning systems, and providing efficient learning support and improvement. Due to this importance, research has been conducted to improve KC models through Difficulty Factors Assessment (DFA) and to evaluate the improvements using the Additive Factors Model (AFM), ultimately creating better learner models (Liu and Koedinger 2017; Koedinger et al. 2010).
@@
Our research proposes a novel approach leveraging the instruction-tuned Large Multimodal Model (LMM) to extract and utilize Knowledge Components (KCs) from educational content. Our method involves parsing educational materials to extract text and images, using the GPT-4o API to identify and describe inherent KCs, and clustering similar components based on sentence embeddings. This automation not only improves KC extraction but also enhances the prediction of student performance on new and unseen content.
@@
Our automatically generated KCs model students’ problem-solving data as effectively as human-created KCs in both KT and AFM.
@@
We publish a reproducible KT benchmark with parsed content data, advancing content-aware Knowledge Tracing methods. 111Codes are available at https://github.com/DoniMoon/LLMKT
Related Works
Utilizing LLMs to improve ITSs and Knowledge Tracing
@@
The emergence of advanced NLP tools in recent years, especially instruction-tuned LLMs like ChatGPT, has significantly enhanced ITSs by providing natural, human-like interactions (Limo et al. 2023; Grassini 2023). Large language models (LLMs) have been shown to provide support in a range of areas, including, but not limited to, planning learning instruction (Hu et al. 2024), scaffolding (Goslen et al. 2024; Liu et al. 2024), and helping students solve math word problems (He-Yueya et al. 2023). They assist learners in multiple ways, such as participating in back-and-forth instructional conversations (Lieb and Goel 2024; Salminen et al. 2024) or providing feedback to students (Dai et al. 2023; Gubelmann et al. 2024).
@@
Moreover, LLMs have also been used to improve the accuracy and performance of the KT models in the backbone of ITSs. LM-KT was proposed to perform KT even when there is no prior problem-solving data from the student. The LM-KT model trains GPT-2 to perform Knowledge Tracing on content without any prior student interaction records (Srivastava and Goodman 2021). In a basic second-language acquisition problem space, LM-KT models student success rates based on the input of natural language sentences as questions. Following LM-KT, other work has shown that leveraging the generalizability of LLMs can enhance the performance of Knowledge Tracing (Neshaei et al. 2024; Zhan et al. 2024) and address cold-start problems (Lee et al. 2024). However, their applicability to other domains is limited, as the language model needs to be fine-tuned on the specific domain.
@@
LLMs have the ability to digest information from source documents which can help with extracting information from large-scale educational data; for example, they have been used for summarize content by intention (Shin et al. 2022) or extracting key points out of educational data (Dagdelen et al. 2024). However, the application of LLMs to improve the performance of KT models, specifically by assisting in the process of KC extraction, has been rare in the literature. We address this gap by providing and evaluating a structured KC extraction method using LLMs in the loop, with the goal of improving the accuracy of KC models while reducing the need for human labeling labor.
Extracting Knowledge Components
@@
Commonly, the assignment of KCs to questions is done by human experts, e.g., instructors (Sun et al. 2019); however, this requires manual labor, making it less suitable for course offerings with a high number of questions. As a result, previous researchers have explored methods to move from expert-annotated KCs towards extracting KCs from the problem information automatically, which can aid in developing more accurate KT models. For example, Thaker, Brusilovsky, and He (2019) proposed a methodology to extract KCs from documents using classical NLP techniques and annotate these documents for application in adaptive online textbooks. Additionally, Sun et al. (2019) validated the accuracy of a method for automatically generating skills for problems by fine-tuning a language model on problem and skill label data to enable computer adaptive test.
@@
Unlike previous studies, we perform the task of tagging the Knowledge Components (KCs) required by questions without using pre-trained problem data or provided schemas of knowledge components. Additionally, we focus on generating entities that clearly correspond to the previous concept of ‘knowledge component,’ rather than simply using terms like skill, knowledge, or tag, and we conduct validation for the KCs. The study most closely related to ours, which analyzes the relation of questions and KCs using LLMs, is Moon et al. (2022). This study uses pre-trained LLMs to verify that if a problem generated from a single KC is a good one, there is a strong dependency on that KC, for the purpose of evaluating automatically generated questions. Conversely, we aim to extract KCs directly from the questions, under the assumption that each question is designed to test specific knowledge.
Methodology
Dataset
@@
We chose to process the OLI datasets from CMU Datashop222https://pslcdatashop.web.cmu.edu/ since the OLI_Statics2011 dataset is a well-known benchmark for knowledge tracing which has publicly available content data(Koedinger et al. 2010; Abdelrahman, Wang, and Nunes 2023). The Open Learning Initiative (OLI) project provides research-based courseware suitable for various class formats and supports advanced research. We gained access to several domains of OLI learning content from CMU Datashop. The preprocessing code for the datasets is publicly available333Currently attached to the submission. The code will be released post anonymous review., and to facilitate reproduction while protecting the content data, the parsed files for each dataset have been uploaded to the corresponding entries in Datashop.
Knowledge Component Extraction
@@ 1. Content Processing
The content data available from the CMU DataShop consists of HTML pages representing the learning materials that students interact with (Koedinger et al. 2010). We parsed this data, extracting text and images from Flash files embedded within the content, and converted MP3 files into text data using the whisper-large-v2 model (Radford et al. 2022). We have processed the image files to request them in the Chat Template from the OpenAI API 444https://platform.openai.com/docs/guides/vision, preserving their positions and order from the original HTML content where the images were located. Each problem was matched with the corresponding steps in the CMU DataShop transaction data, creating datasets for five subjects: Statics, Psychology, Biology, Computing, and French.
2. Knowledge Component Extraction
@@
Using the OpenAI API for the GPT-4o model, we extracted the knowledge components for each problem as shown in Fig 2. Each problem can have multiple knowledge components, with each component consisting of a name field (1-3 words) and a description field (1-2 sentences).
3. Clustering Knowledge Components
@@
The extracted KCs consisted of natural language sentences, which occasionally referred to the same topic but with different sets of words. To utilize these components for problem correlation and be able to assign the same identifiers to semantically similar KCs, we needed to group them together. To do this, we computed sentence embeddings for each component and performed clustering based on similarity. The optimal number of clusters was determined by maximizing the silhouette score of the clustering (Rousseeuw 1987). We compared the performance of the Sentence-T5-XXL model (Ni et al. 2022) and OpenAI’s text-embedding-3-large model for this task 555https://openai.com/index/new-embedding-models-and-api-updates/.

### v2

@@ Abstract
Knowledge tracing models have enabled a range of intelligent tutoring systems to provide feedback to students. However, existing methods for knowledge tracing in learning sciences are predominantly reliant on statistical data and instructor-defined knowledge components, making it challenging to integrate AI-generated educational content with traditional established methods. We propose a method for automatically extracting knowledge components from educational content using instruction-tuned large multimodal models. We validate this approach by comprehensively evaluating it against knowledge tracing benchmarks in five domains. Our results indicate that the automatically extracted knowledge components can effectively replace human-tagged labels, offering a promising direction for enhancing intelligent tutoring systems in limited-data scenarios, achieving more explainable assessments in educational settings, and laying the groundwork for automated assessment. 111Codes are available at https://github.com/DoniMoon/LLMKT
keywords:
Knowledge Tracing, Multimodal Models, Multimedia Question Information, Knowledge Components1 Introduction
A flowchart illustrating the four main steps of the experiment. From top to bottom: “Problem Content” feeds into “LMM Inference,”
“Knowledge Descriptions” feeds into “Calculate Sentence Embedding,”
“Embedding Vectors” feeds into “Clustering,” and
“Knowledge Components” feed into “Evaluate.” The final step, “Evaluate,” references two validation tasks: Additive Factor Model (AFM) and Knowledge Tracing (KT), which are linked to Tables 3 and 4 in the paper.
@@
Intelligent Tutoring Systems (ITS) are advanced computer programs that provide personalized and adaptive educational instruction to learners. For over half a century, they have been a subject of active research and discussion in the interdisciplinary field of education and artificial intelligence [28]. These systems integrate techniques from artificial intelligence to deliver tailored instruction, dynamically adjusting to the needs and progress of individual students. By simulating one-on-one tutoring, ITS offer immediate and specific feedback, enhancing student engagement and improving learning outcomes. This combination of real-time adaptability and personalized feedback makes ITS a valuable tool in modern education, bridging the gap between traditional classroom methods and individualized learning approaches [31].
@@
Knowledge Tracing (KT) is a foundational task in Intelligent Tutoring Systems (ITS), aiming to model a student’s knowledge state and predict future performance on educational tasks. Traditional KT models rely heavily on statistical techniques to analyze historical problem-solving data. Early approaches, such as Bayesian Knowledge Tracing (BKT), used hidden Markov models to estimate the probability of a student knowing a particular skill at any given time [7].
@@
More recent models have leveraged advancements in machine learning, particularly deep learning, to enhance predictive accuracy [1]. Deep Knowledge Tracing (DKT) was one of the first models to apply recurrent neural networks (RNNs) to KT, demonstrating significant improvements over BKT by capturing the sequential nature of students’ learning processes. Subsequent models such as Self-Attentive Knowledge Tracing (SAKT) and Separated Self-AttentIve Neural Knowledge Tracing (SAINT) have further refined these approaches [5, 32].
@@
Knowledge Components (KCs) are defined within the Knowledge-Learning-Instruction framework as acquired units of cognitive function or structure that can be inferred from performance on a set of related tasks [19]. While we cannot directly observe the changes in a student’s KCs, we infer them through interactions during assessment and instruction. In Knowledge Tracing, each problem is often labeled with its corresponding KCs, which are typically assigned by human experts based on their expected relevance to the problem and statistically validated for their ability to explain student performance.
@@
There are several methods to define these KCs. [35] introduces four categories of domain modeling: ‘KCs as disjoint sets’, ‘Multiple KCs per item’, ‘Hierarchy of KCs’, and ‘KCs with prerequisites’. In the KT domain, the most common approaches are modeling KCs as disjoint sets or mapping multiple KCs to an item. The simplest approach is to map each question to a single KC, as used in Bayesian KT, SAKT, and SAINT. When mapping multiple KCs, the relationship between KCs and items is represented in a Q-matrix, where rows represent questions and columns represent KCs. Cognitive Diagnosis Models (CDM) like DINA, NIDA, and generalized DINA use this Q-matrix for KT [10], thus supporting multiple KCs labeled for each question.
A schematic flow for extracting and grouping knowledge components from a sample question about force direction on a cable. On the left, a monitor labeled “GPT-4o” displays the question “What is the direction and sense of the force exerted on the displayed section of the cable?” Below it, bullet points summarize extracted knowledge components (e.g., identifying force directions, understanding tension in cables, and analyzing pulleys). On the right, these components are grouped under two headings: “Moment calculation” (focusing on coordinate points and torque) and “Trigonometric functions” (focusing on formulas for tension). Colored stacks of bars illustrate clustering of the knowledge components based on sentence embeddings, forming semantically related groups.
@@
Well-defined KCs are essential for accurately predicting learner performance, building effective adaptive learning systems, and providing efficient learning support and improvement. Due to this importance, research has been conducted to improve KC models through Difficulty Factors Assessment (DFA) and to evaluate the improvements using the Additive Factors Model (AFM), ultimately creating better learner models [23, 18].
@@
Our research proposes a novel approach leveraging the instruction-tuned Large Multimodal Model (LMM) to extract and utilize Knowledge Components (KCs) from educational content. Figure 1 illustrates the overall architecture of our approach. Our method involves parsing educational materials to extract text and images, using the GPT-4o API to identify and describe inherent KCs, and clustering similar components based on sentence embeddings. This automation not only improves KC extraction but also enhances the prediction of student performance on new and unseen content.
@@
Our automatically generated KCs model students’ prob-lem-solving data as effectively as human-created KCs in both KT and AFM.
@@
We publish a reproducible KT benchmark with parsed content data, advancing content-aware Knowledge Tracing methods.
2 Related Work
2.1 Utilizing LLMs to improve ITS and Knowledge Tracing
@@
The emergence of advanced NLP tools in recent years, especially instruction-tuned LLMs like ChatGPT, has significantly enhanced ITS by providing natural, human-like interactions [22, 13]. Large language models (LLMs) have been shown to provide support in a range of areas, including, but not limited to, planning learning instruction [16], scaffolding [12, 25], and helping students solve math word problems [15]. They assist learners in multiple ways, such as participating in back-and-forth instructional conversations [21, 38] or providing feedback to students [9, 14].
@@
Moreover, LLMs have also been used to improve the accuracy and performance of the KT models in the backbone of ITS. LM-KT was proposed to perform KT even when there is no prior problem-solving data from the student. The LM-KT model trains GPT-2 to perform Knowledge Tracing on content without any prior student interaction records [41]. In a basic second-language acquisition problem space, LM-KT models student success rates based on the input of natural language sentences as questions. Following LM-KT, other work has shown that leveraging the generalizability of LLMs can enhance the performance of Knowledge Tracing [29, 45] and address cold-start problems [20]. However, their applicability to other domains is limited, as the language model needs to be fine-tuned on the specific domain.
@@
LLMs have the ability to digest information from source documents which can help with extracting information from large-scale educational data; for example, they have been used for summarize content by intention [40] or extracting key points out of educational data [8]. However, the application of LLMs to improve the performance of KT models, specifically by assisting in the process of KC extraction, has been rare in the literature. We address this gap by providing and evaluating a structured KC extraction method using LLMs in the loop, with the goal of improving the accuracy of KC models while reducing the need for human labeling labor.
2.2 Extracting Knowledge Components
@@
Commonly, the assignment of KCs to questions is done by human experts, e.g., instructors [42]; however, this requires manual labor, making it less suitable for course offerings with a high number of questions. As a result, previous researchers have explored methods to move from expert-annotated KCs towards extracting KCs from the problem information automatically, which can aid in developing more accurate KT models. For example, a methodology to extract KCs from documents using classical NLP techniques and annotate these documents for application in adaptive online textbooks has been proposed [43]. Additionally, there is a study that validated the accuracy of a method for automatically generating skills for problems by fine-tuning a language model on problem and skill label data to enable computer adaptive tests [42].
@@
Unlike previous studies, we perform the task of tagging the Knowledge Components (KCs) required by questions without using pre-trained problem data or provided schemas of knowledge components. Additionally, we focus on generating entities that clearly correspond to the previous concept of ‘knowledge component’, rather than simply using terms like skill, knowledge, or tag, and we conduct validation for the KCs. The study most closely related to ours analyzed the relation of questions and KCs using LLMs [27]. This study uses pre-trained LLMs to verify that if a problem generated from a single KC is a good one, there is a strong dependency on that KC, for the purpose of evaluating automatically generated questions. Conversely, we aim to extract KCs directly from the questions, under the assumption that each question is designed to test specific knowledge.
3 Methodology
As illustrated in Figure 2, our overall pipeline involves preprocessing student interaction data enriched with content information, extracting KCs, evaluating the quality of the extracted KCs, and analyzing their utility across various KT methods.
Two line charts display clustering metrics for varying numbers of clusters (from 0 to 200) in the “oli_statics” dataset. The left chart shows the Elbow Method, with the y-axis labeled “WCSS,” which decreases as the number of clusters increases, forming a downward-sloping curve. The right chart shows the Silhouette Score, which rises gradually on the y-axis as the number of clusters increases, indicating how well-separated the clusters are.
3.1 Dataset
@@
We chose to process the OLI datasets from CMU Datashop222https://pslcdatashop.web.cmu.edu/ since the OLI_Statics2011 dataset is a well-known benchmark for knowledge tracing which has publicly available content data[18, 1]. The Open Learning Initiative (OLI) project provides research-based courseware suitable for various class formats and supports advanced research. We gained access to several domains of OLI learning content from CMU Datashop. The preprocessing code for the datasets is publicly available, and to facilitate reproduction while protecting the content data, the parsed files for each dataset have been uploaded to the corresponding entries in Datashop.
From CMU Datashop, all transaction data was extracted by selecting “all data” in the “Export” tab and choosing “By Transaction” in the detailed options. The content data was collectively downloaded by clicking the “Download Problem Content” button in the “Dataset Info/Problem List” tab. We used the swftools package 333http://www.swftools.org/ in the Windows OS environment to extract images from the SWF file, which is not supported anymore due to the deprecation of Flash support.
all mp3 files are converted into text files, using whisper-large-v2 [36].
3.1.1 OLI Engineering Statics
For the Statics dataset, we used the Fall 2011 version444https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=507. While other versions included a wider variety of KC models, the number of users and transactions was smaller compared to other statics benchmarks. Therefore, we considered the Fall 2011 version to be the superset and used it. The KC model named F2011 was used as the human tag. The original 361,092 transactions were reduced by about half to 189,047.
3.1.2 OLI Principles of Computing
The content of the OLI computing dataset 555https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=1806 was different from other subjects, as it contained content from multiple subjects under the root directory. To maintain consistency with the method used for processing other datasets, only the content from Principles_of_Computing_v_1_10, which includes the version in the subject name, was used. The KC models principles_of_computing_1_13 and principles_of_computing_1_10 were used to calculate AFM scores. The transaction data was filtered to include only those containing this content. As a result, the original transaction dataset, which had 37,233 rows, was reduced to 16,951 rows.
3.1.3 OLI French
We used the ‘French1 - Spring 2014’ dataset for the French dataset666https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=918. Since it included various KC models, we used Bonnie’s Model, which had the best AFM performance. To train AFM, Level4 KC model, which also reports good AFM performance, is used. This dataset contained many items with voice mp3 files, so we converted these files to text using the whisper-large-v2 model and inserted the transcriptions into the question text. At the position in the HTML where the mp3 execution was embedded, we prefixed the converted text with ‘[transcription of embedded mp3 file]:’. Out of the total 278,489 rows of transactions, the final remaining data consisted of 53,255 rows.
3.1.4 OLI Biology
We used the ‘Oli_biology’ dataset from the ‘Bridge to Success’ project for the Biology data777https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=1148. The KC model named intro_biology-1_0 was used as the human label. After processing, 3,285,685 rows remained out of the original 5,852,795 rows of transactions.
3.1.5 OLI Psychology
We used the ‘Psychology MOOC GT - Spring 2013’ dataset for the Psychology data888https://pslcdatashop.web.cmu.edu/DatasetInfo?datasetId=863 and utilized the KC model named psychology-1-4. After processing, 1,935,496 rows remained out of the original 2,493,609 rows of transactions.
Two boxes list example knowledge components from a statics dataset. The upper box, labeled “Trigonometric Relationships,” contains items about decomposing forces into perpendicular components and applying trigonometric functions to resolve vectors. The lower box, labeled “Moment Calculation,” includes items about determining force direction and sense, recognizing how forces contribute to equilibrium, and computing the moment of a force. Each bullet point corresponds to a specific sub-topic within these clusters.
3.2 Knowledge Component Extraction
@@ 3.2.1 Content Processing
The content data available from the CMU DataShop consists of HTML pages representing the learning materials that students interact with [18]. We parsed this data, extracting text and images from Flash files embedded within the content, and converted MP3 files into text data using the whisper-large-v2 model [36]. We processed the image files to embed them in the Chat Template for the OpenAI API 999https://platform.openai.com/docs/guides/vision, preserving their positions and order from the original HTML content where the images were located. Each problem was matched with the corresponding steps in the CMU DataShop transaction data, creating datasets for five subjects: Statics, Psychology, Biology, Computing, and French.
3.2.2 Knowledge Component Extraction Prompt
@@
Using the OpenAI API for the GPT-4o model, we extracted the knowledge components for each problem as shown in Figure 2. Each problem can have multiple knowledge components, with each component consisting of a name field (1–3 words) and a description field (1–2 sentences). We applied this basic zero-shot prompt from Fig. 2 only once and, being satisfied with the qualitative output, did not undertake any further prompt tuning.
3.2.3 Clustering Knowledge Components
@@
The extracted KCs consisted of natural language sentences, which occasionally referred to the same topic but with different sets of words. To utilize these components for problem correlation and be able to assign the same identifiers to semantically similar KCs, we needed to group them together. To do this, we computed sentence embeddings for each component and performed clustering based on similarity. The optimal number of clusters was determined by maximizing the silhouette score of the clustering [37]. We compared the performance of the Sentence-T5-XXL model [30] and OpenAI’s text-embedding-3-large model for this task 101010https://openai.com/index/new-embedding-models-and-api-updates/.
Figure 3 shows the WCSS and Silhouette Score of the K-means clustering method as the number of clusters varies from 2 to 200 in the oli_statics dataset. Due to the instability of the Silhouette Score when the number of clusters is very small, we analyzed cases where the number of clusters is greater than 10. Each point in Figure 6 represents the local maximum Silhouette Score within one of the 10 bins divided between the range of 10 to 200 clusters, and the AFM performance was measured using these cluster numbers.
Meanwhile, for zero-shot implementation, using more than 100 KCs, which maximizes the Silhouette Score, made it impossible to have at least one problem with each KC in both the train and test splits. Therefore, we used the number of clusters at the local maximum of the third bin, where such splits were feasible for all datasets. The selected numbers of clusters were 52, 63, 52, 61, and 49 for computing, statics, French, psychology, and biology, respectively.
3.3 Knowledge Component Quality Evaluation
The KCs used in this validation were evaluated by measuring the silhouette score of K-means clustering from 2 to 200 clusters. We automatically determined the optimal number of clusters by detecting the knee point of the silhouette score change [37, 39]. To further verify the quality of KCs based on the level of clustering, we divided the entire range of 2 to 200 clusters into ten sections. For each section, we identified the point where the silhouette score was at its local maximum and observed the performance change of the AFM when KCs were generated with the corresponding number of clusters.
RMSE was measured not only for the entire dataset but also in environments where student ID information and item ID information were masked, respectively. This ablation experiment was done to identify which biases the model was utilized to demonstrate performance. For performance measurement, we utilized PyAFM, a Python implementation of the AFM [26].
To better understand the Knowledge Components (KCs) categorized through clustering, the most frequent name within each cluster was selected as the representative name for that cluster, alongside the descriptions included in that cluster. As illustrated in Figure 4, these names typically consist of 2-3 words, such as Trigonometric Relationships or Moment Calculation. In most clusters, approximately one-third of the items shared the same name, which was then chosen as the mode and designated as the cluster’s representative name.
3.4 Knowledge Tracing Baselines Using KCs
Using the extracted knowledge components, we used code from prior work [11] to conduct KT with IRT, PFA, DAS3H, SAKT, and DKT. In our implementation of PFA, we also considered incorporating a student-specific intercept—an enhancement reported to boost performance in recent work [6]—but observed degraded results across all configurations; consequently, we omitted student intercept encoding in our final PFA setup.
In the experiments of the previous section, the performance of the OpenAI embedding model was superior to that of the T5-XXL model, so we evaluated the KCs generated using the OpenAI text-embedding-3-large model. Among these, IRT is an algorithm that does not use KCs, while PFA and DAS3H use a Q-matrix. SAKT and DKT operate in environments with disjoint KCs, so we used joint skill assignments. Joint skill considers the combination of KCs assigned to each question as a single KC [44]. For example, a problem with KC 1 and KC 2 is treated as having a distinct KC different from a problem with only KC 1.
3.5 Detailed Experiment Settings
3.5.1 LMM Inference
A rectangular prompt box labeled “Prompt” displays instructions for extracting knowledge components needed to solve a question. It specifies that each component should have a short name (2–4 words) and a single-sentence description, and the output should be in JSON format. An example JSON snippet is shown, containing one knowledge component with a name and description.
[…]
3.5.2 Zero-Shot Knowledge Tracing
We implemented zero-shot knowledge tracing using the same codebase111111https://github.com/theophilegervet/learner-performance-prediction that was used for KT implementation[11]. The logistic regression setup was configured to experiment with various combinations of features, specifically utilizing the s, sc, tc, tw, w, and a tags. These options allow us to use features that record which KCs are present, how many times each KC appears in the user’s history, the total number of problems the user has solved, the total number of problems the user has correctly solved, and how many attempts were made for each problem within a time window. When selecting these options, we understood the code and chose all relevant options, but did not compare with other options due to the absence of a validation set.
3.6 Quality of the Generated Knowledge Components
| Generated Name | Description | KC Name |
| Item ID: 1273 | ||
| Reading comprehension | Understand the text of the question and the options provided. | identifying correct option |
| Multiple choice format | Recognize the structure of a multiple choice question and how to select an answer. | understanding question format |
| Decision making | Decide between the given options based on the question’s requirements. | identifying correct option |
| Item ID: 1097 | ||
| Summation of Forces | Understand that F y = 0 denotes the summation of all forces in the y-direction equaling zero. | Force Equilibrium |
| Force Equilibrium | Recognize that the condition F y = 0 implies a state of force equilibrium in the vertical direction. | Force Equilibrium |
| Multiple Choice Questions | Know how to interpret and answer multiple-choice questions. | multiple choice format |
| Selecting Correct Answer | Identify the correct option based on given conditions and context. | identifying correct option |
| Generated Name | Description | KC Name |
| Item ID: 2066 | ||
| Human Eye Anatomy | Understanding the different parts of the human eye and their functions. | Young-Helmholtz theory |
| Iris Location | The iris is positioned between the cornea and lens, controlling the amount of light that enters the eye. | Young-Helmholtz theory |
| Item ID: 709 | ||
| Reasonable Mind Concept | Understanding that the decision involves logical planning and time management. | True or False Questions |
| Wise Mind Concept | Recognizing the balance between emotional and reasonable mindsets, though not applicable here. | Identifying emotions |
| Emotional Mind Concept | Understanding that decisions driven purely by emotions are not being considered in this scenario. | Identifying emotions |
| Time Management Skills | The ability to plan and allocate time effectively for various activities including work, study, and social events. | True or False Questions |
Three side-by-side line charts show how AFM performance (on the y-axis, labeled RMSE) varies with the number of clusters (x-axis from 0 to 200). Each chart represents a different method: Stratified CV (left), Student CV (middle), and Item CV (right). In each chart, there are three lines for the different subject areas: Computing (blue), French (orange), and Statics (red). The trends illustrate how RMSE changes as the number of clusters increases for each subject and cross-validation approach.
3.6.1 AFM Analysis
For AFM inference, we used the Python implementation compatible with the CMU DataShop format121212https://github.com/cmaclell/pyAFM. The mapped KCs were post-processed in the same scheme as when transactions are exported with the ‘By Transaction’ option from CMU DataShop, and the inference was performed using this code. We were unable to conduct AFM analysis on the Psychology and Biology datasets because the memory usage of the code increases proportionally with the transaction length. The required memory exceeded 40GB, which could not be handled by the computational resources used in this study, specifically a MacBook M3 Pro with 36GB of shared memory. We used 3-fold cross-validation, adopting the default hyperparameters from the original AFM implementation [26].
4 Results